# Learning Value of Information towards Joint Communication and Control in 6G V2X

### v1

@@ I Introduction
Connected Autonomous Vehicles (CAVs) integrate automated control with advanced communication technologies to achieve self-driving capabilities with minimal or no human intervention. Cellular Vehicle-to-Everything (C-V2X) technologies are advancing rapidly and are on the brink of commercial deployment. As C-V2X evolves towards six-generation (6G), CAVs are anticipated to be a key application in the 6G era [1].
@@
Unlike Autonomous Vehicles (AVs) without communication capabilities, CAVs can achieve a more accurate and comprehensive perception of the driving environment by leveraging critical information exchanged through Vehicle-to-Everything (V2X) communications, which is referred to as ‘V2X information’ in this paper. This enables them to make more reliable and efficient driving decisions by integrating V2X information with local sensory data [2].
@@
Recent years have witnessed a growing interest in applying data-driven Machine Learning (ML) to CAVs. Deep Reinforcement Learning (DRL), which combines deep learning (DL) with reinforcement learning (RL), has shown significant advantages over traditional model-driven control and optimization methods, as it requires fewer restrictive assumptions about the stochastic properties of system dynamics [3, 4, 5, 6]. Leveraging DRL for both vehicle control and V2X communication, an intelligent, integrated decision-making system presents a promising solution for ensuring safe and efficient autonomous driving (AD). To enable this, Markov Decision Process (MDP) models must be conceived for framing the vehicle control and V2X communication problems, respectively. The system would be comprised of two types of agents - a control agent and a communication agent - that employ DRL algorithms to solve the MDP models and learn optimal policies for decision-making. Notably, the control and communication decisions may operate at different time scales [7].
@@
The control agent makes decisions on vehicle control inputs based on available sensory and V2X information. These inputs are then executed by actuators, such as steering, throttle, brakes, and gear shifting, enabling autonomous vehicle operation. It is widely recognized that providing the control agent with more information at the time of decision-making can lead to better decisions. However, an overwhelming amount of information can be toxic to the control agent and greatly increases the computational complexity of making decisions due to the curse of dimensionality problem [8]. Ideally, the control agent would have access to the necessary and sufficient information to make optimal decisions. In practice, however, information comes at a cost and can be imperfect. This paper mainly focuses on the communication costs incurred for information dissemination and the imperfections in information resulting from non-ideal communication.
@@
Since the information is utilized by control agents for decision-making, the VoI should be evaluated by these agents. The communication agents - whether at the base station (BS), roadside infrastructure, CAVs and other information sources - can leverage the estimated VoI, alongside traditional communications performance metrics such as throughput and delay, to optimize their communication decisions.
The VoI depends on how the control agent reacts to missing or imperfect information by adjusting its decisions, which in turn relies on the type of controllers used. Traditionally, AV controllers have been designed based on classical control theory, such as linear controller, $\mathcal{H}_{\infty}$ controller, and Sliding Mode Controller (SMC) [11]. In this paper, we focus on VoI when DRL-based AV controllers are employed. Our results, however, extend to any AV controller designed by solving MDPs, including both model-based dynamic programming/optimal control and model-free reinforcement learning methods.
@@ I-B1 Related Works
Information theory provides powerful tools for quantifying the amount of uncertainty in the value of a random variable, and has had a profound influence on the design of communication systems. However, as Howard indicates in his seminal work, “no theory that involves just the probabilities of outcomes without considering their consequences could possibly be adequate in describing the importance of uncertainty to a decision maker” [9]. Howard introduces the theory of VoI, and asserts that numerical values can be assigned to the elimination or reduction of any uncertainty. To explain the concept and intuition of VoI, Howard uses a bidding problem as an example, where the expected profit of eliminating any certainty is used as the VoI. However, Howard’ work does not provide a rigorous definition or mathematical formula for calculating VoI. Following his work, various definitions of VoI have been proposed for different applications and objectives. A survey on VoI in Wireless Sensor Network (WSN) and Internet of Things (IoT) by Alawad et al. [10] provides a structured literature review on the fragmented works in this area. Despite these contributions, a theoretical foundation for VoI is still lacking, which would allow VoI to be estimated using uniform formulas for different applications.
@@ I-B2 Contribution 2
To derive the information theory-based VoI, we calculate the conditional KL divergence between the probability distribution given by the product of the transition models of the original state and the V2X information, and the probability distribution given by the transition model of the augmented state, which inclues both the original state and the V2X information. While information theory-based VoI is easier to derive given the model of the environment, utility-based VoI has several important advantages. Therefore, we introduce a DRL-based method to estimate the utility-based VoI by offline training policies with missing/imperfect and full/perfect information, respectively. Then, the regret and advantage function can be estimated to quantify the EVoI and IVoI, respectively.
@@ I-C1 Related Works
In V2X communication, there is a growing interest aimed at designing RRM mechanisms to minimize the Age of Information (AoI) [18, 19, 20, 21], rather than focusing solely on traditional metrics such as throughput and delay. While AoI measures the importance of information by assessing its timeliness, it does not account for the broader benefits that information can provide. In contrast, VoI directly quantifies the advantage of information in enhancing the performance of specific tasks. Unfortunately, there is currently a lack of studies focused on VoI-aware RRM.
In MDP and RL communities, the problems of “What to Communicate” and “When to Communicate” have been studied in the context of Decentralized-POMDP (Dec-POMDP) [22, 23] and Multi-agent RL (MARL) [24, 25]. The goal in these studies is to simultaneously learn both a communication policy and an acting policy to maximize performance while adhering to communication constraints. However, VoI is rarely used in this line of research [23]. One negative consequence of this neglect is that it is challenging to explicitly express the trade-off between maximizing the performance of considered task and that of traditional communication services. As a result, communication constraints in these studies are often simplified, typically involving restrictions on the number of messages or the agents allowed to transmit.
@@ II System Model
Figure 1: Illustration of system model
@@
Figure 2: Examples of typical scenarios
@@
where $\rho_{i}$ is a time constant representing the driveline dynamics of vehicle $i$.
We denote the headway of vehicle $i$ at time $t$, i.e., bumper-to-bumper distance between $i$ and its preceding vehicle (predecessor) $i-1$, by $d_{i}$, which satisfies
@@
$d_{c,i}=c_{i}+h_{i}v_{i},$ (3)
where $c_{i}$ is a constant standstill distance for vehicle $i$ and $h_{i}$ is the desired time-gap of vehicle $i$.
@@
Let $x_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$, where the subscript $k$ indicates the values of the variables at the beginning of control interval $k$. The system dynamics in discrete time are derived from (1) on the basis of forward Euler discretization:
@@
$A_{i}=\begin{bmatrix}1&T&-h_{i}T\\ 0&1&-T\\ 0&0&1-\frac{T}{\tau_{i}}\end{bmatrix},B_{i}=\begin{bmatrix}0\\ 0\\ \frac{T}{\tau_{i}}\end{bmatrix},C_{i}=\begin{bmatrix}0\\ T\\ 0\end{bmatrix}.$ (6)
At the beginning of each control interval, the control agent of vehicle $i$ has to determine the control input $u_{i,k}$ based on the observation of the state. The velocity $v_{i,k}$ and acceleration $acc_{i,k}$ can be measured locally, while the control errors $e_{pi,k}$ and $e_{vi,k}$ can be quantified by a radar unit mounted at the front of the vehicle. On the other hand, vehicle $i$ can only obtain the acceleration $acc_{i-1,k}$ of the predecessor through V2X communications. To determine the optimal action $u_{i,k}$ for $k=0,1,\cdots,K-1$, we must decide what, when, and how to share information by V2X communications from predecessor $i-1$ to its follower $i$.
@@ IV-A SSDP Basics
    which is derived from the predecessor $i-1$’s dynamic model $\dot{acc}_{i-1}=-\frac{1}{\tau_{i-1}}acc_{i-1}+\frac{1}{\tau_{i-1}}u_{i-1}$ based on forward Euler discretization;
@@
    $r(S_{i,k},a_{i,k})=-\{(e_{pi,k})^{2}+\alpha(e_{vi,k})^{2}+\beta(a_{i,k})^{2}\},$ (17)
    where $\alpha$ and $\beta$ are the weights that are positive and can be adjusted to determine the relative importance of minimizing the position error, velocity error and the control input.
@@ V-A Utility-based VoI
(Value of Missing Information (VoMI)) The VoMI is defined as the difference in the performance of the optimal policy $\pi^{*}$ for the SSDP, where the state $S_{k}$ does not include information $I_{k}$, and the performance of the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes $I_{k}$.
@@
(Value of Imperfect Information (VoII)) The VoII is defined as the difference in the performance of the optimal policy $\tilde{\pi}^{\rm ip*}$ for the augmented-state SSDP with imperfect information, where the state $\tilde{S}_{k}^{\mathrm{ip}}$ is constructed based on the imperfect information $I_{k}^{\rm ip}$, and the performance of the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP $\tilde{S}_{k}=\{S_{k},I_{k}\}$, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes the perfect information $I_{k}$.
Note that both VoMI and VoII correspond to the performance difference between a policy with inferior performance and a policy with superior performance. The former policy is the optimal policy based on partial/imperfect observation (state with partial/imperfect information), while the latter policy is the optimal policy based on full/perfect observation (state with full/perfect information). Depending on how the performance difference is measured, both VoMI and VoII can be further divided into two categories: Expected Cumulative VoI and Immediate VoI, which are discussed in the next subsection.
@@
For ease of understanding, we use unified notations $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}$ to represent the inferior policy and superior policy in the VoMI and VoII definitions, respectively. Additionally, $S_{k}^{\rm{inf}}$ and $S_{k}^{\rm{sup}}$ are used to represent the partial/imperfect observation and full/perfect observation, respectively.
@@
(Expected Cumulative VoI (EVoI)) The EVoI $\bar{\Xi}$ is defined as the difference in the performance of the optimal policy $\pi^{\rm{inf}}$ based on the partial/imperfect observation and the optimal policy $\pi^{\rm{sup}}$ based on the full/perfect observation, respectively. The performance difference is quantified as
@@
The EVoI indicates the performance loss as a consequence of the control agent consistently receiving the partial/imperfect observations at every control interval. In the terminology of RL, EVoI corresponds to the regret of inferior policy $\pi^{\rm{inf}}$. It is worth noting that the EVoI is a non-positive value, with a larger EVoI indicating less performance loss. When the EVoI reaches its maximum value of zero, no performance loss is experienced.
@@
(Immediate Value of Information (IVoI)) Given the full/perfect observation $S_{k}^{\rm{sup}}$ and partial/imperfect observation $S_{k}^{\rm{inf}}$ at control interval $k$, the IVoI $\xi_{k}$ is defined as the difference in the performance starting from control interval $k$, where the decision at control interval $k$ is made by the optimal policy $\pi^{\rm{inf}}$ based on the partial/imperfect observation and the optimal policy $\pi^{\rm{sup}}$ based on the full/perfect observation, respectively, while the rest of the sequential control decisions are made by $\pi^{\rm{sup}}$ based on the full/perfect observations $S_{k+1}^{\rm{sup}},S_{k+2}^{\rm{sup}},\dots$, i.e.,
@@
where $a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})$ indicates that the action $a_{k}^{\rm{inf}}$ is sampled according to $\pi^{\rm{inf}}$.
The IVoI indicates the performance loss as a consequence of the control agent receiving the partial/imperfect observations only at a specific control interval $k$, while consistently obtaining full/perfect observations throughout the remaining time. Moreover, IVoI is calculated for a specific full/perfect and partial/imperfect observation pair at control interval $k$. In other words, IVoI reveals that if the full/perfect observation at control interval $k$ is $S_{k}^{\rm{sup}}=s$, what is the performance loss if the agent only has access to the partial/imperfect observation $S_{k}^{\rm{inf}}$ instead? In the terminology of RL, IVoI corresponds to the advantage function of the optimal policy $\pi^{\rm{sup}}$. Similar to EVoI, IVoI is also a non-positive value, where maximizing IVoI leads to minimizing the performance loss.
@@
Therefore, utility-based VoI can be categorized into two main types: EVoI and IVoI. Each of these types is further divided into two subcategories: VoMI and VoII. As illustrated in Table. I, this classification results in four distinct categories of utility-based VoI: Expected VoMI (EVoMI), Expected VoII (EVoII), Immediate VoMI (IVoMI), and Immediate VoII (IVoII). According to Definition 5 and 6, the policy/state/action $\pi^{\rm{inf}}$/$S_{k}^{\rm{inf}}$/$a_{k}^{\rm{inf}}$ in (22) and (23) is replaced by $\pi^{*}$/$S_{k}$/$a_{k}$ and $\tilde{\pi}^{\rm ip*}$/$\tilde{S}_{k}^{\rm ip}$/$a_{k}^{\rm ip}$ for VoMI and VoII, respectively, while $\pi^{\rm{sup}}$/$S_{k}^{\rm{sup}}$ in (22) and (23) is replaced by $\tilde{\pi}^{*}$/$\tilde{S}_{k}$.
VoMI VoII
EVoI EVoMI EVoII
$\bar{\Xi}=J_{\pi^{*}}-J_{\tilde{\pi}^{*}}$ $\bar{\Xi}=J_{\tilde{\pi}^{\rm ip*}}-J_{\tilde{\pi}^{*}}$
IVoI IVoMI IVoII
$\xi_{k}=A_{\tilde{\pi}^{*}}(\tilde{S}_{k},a_{k})|_{a_{k}\sim\pi^{*}(\cdotp|S_{k})}$ $\xi_{k}=A_{\tilde{\pi}^{*}}(\tilde{S}_{k},a_{k}^{\rm ip})|_{a_{k}^{\rm ip}\sim\tilde{\pi}^{\rm ip*}(\cdotp|\tilde{S}_{k}^{\rm ip})}$
TABLE I: Four categories of utility-based VoI

### v2

@@ I Introduction
Connected Autonomous Vehicles (CAVs) integrate automated control with advanced communication technologies to achieve self-driving capabilities with minimal or no human intervention. Cellular Vehicle-to-Everything (C-V2X) technologies are advancing rapidly and are on the brink of commercial deployment. Unlike Autonomous Vehicles (AVs) without communication capabilities, CAVs can achieve a more accurate and comprehensive perception of the driving environment by leveraging critical information exchanged through Vehicle-to-Everything (V2X) communications, which is referred to as ‘V2X information’ in this paper. This enables them to make more reliable and efficient driving decisions by integrating V2X information with local sensory data [1, 2, 3].
As C-V2X evolves towards sixth-generation (6G) networks to support CAVs, several key challenges emerge, including Artificial Intelligent (AI)-native architectures, integrated sensing and communication, and increasingly stringent Ultra-Reliable Low-Latency Communication (URLLC) requirements[4]. Among these, a fundamental requirement is the timely and efficient utilization of limited communication resources to transmit critical information for autonomous driving. Unlike traditional networks that focus primarily on raw data transmission, 6G envisions a tight integration of communication and control functions in CAVs, enabling goal-oriented communication that prioritizes task-relevant information. This paradigm shift demands communication systems that are not only fast and reliable but also control-aware, capable of identifying and delivering critical information essential for real-time driving decisions.
@@
Recent years have witnessed a growing interest in applying data-driven Machine Learning (ML) to CAVs. Deep Reinforcement Learning (DRL), which combines deep learning (DL) with reinforcement learning (RL), has shown significant advantages over traditional model-driven control and optimization methods, as it requires fewer restrictive assumptions about the stochastic properties of system dynamics [5, 6, 7, 8, 9]. Leveraging DRL for both vehicle control and V2X communication, an intelligent, integrated decision-making system presents a promising solution for ensuring safe and efficient autonomous driving (AD). To enable this, Markov Decision Process (MDP) models must be conceived for framing the autonomous driving (AD) and V2X communication problems, respectively. The system would be comprised of two types of agents - a control agent and a communication agent - that employ DRL algorithms to solve the MDP models and learn optimal policies for decision-making. Notably, the control and communication decisions may operate at different time scales [10].
@@
The control agent makes AD decisions based on available sensory and V2X information. These decisions can occur at various levels, including perception understanding, behavior and motion planning, vehicle control, or an end-to-end approach that encompasses all levels [11]. In this paper, we focus on vehicle control decisions to illustrate our proposed methodology, although it is applicable to all types of AD decisions. The vehicle control decisions made by the control agent — referred to as control inputs — are executed by actuators such as the steering system, throttle, brakes, and gear shifter, thereby enabling autonomous vehicle operation. It is widely recognized that providing the control agent with more information at the time of decision-making can lead to better decisions. However, an overwhelming amount of information can be toxic to the control agent and greatly increases the computational complexity of making decisions due to the curse of dimensionality problem [12]. Ideally, the control agent would have access to the necessary and sufficient information to make optimal decisions. In practice, however, information comes at a cost and can be imperfect. This paper mainly focuses on the communication costs incurred for information dissemination and the imperfections in information resulting from non-ideal communication.
@@
Since the information is utilized by control agents for decision-making, the VoI should be evaluated by these agents. The communication agents - whether at the base station (BS), roadside infrastructure, CAVs and other information sources - can leverage the estimated VoI, alongside traditional communications performance metrics such as throughput and delay, to optimize their communication decisions [15].
The VoI depends on how the control agent reacts to missing or imperfect information by adjusting its decisions, which in turn relies on the type of control algorithm employed. Traditionally, AV control algorithms have been designed based on classical control theory, such as linear controller, $\mathcal{H}_{\infty}$ controller, and Sliding Mode Controller (SMC) [16]. In this paper, we focus on VoI when DRL-based AV control algorithms are employed. However, our results are not limited to DRL and can be extended to any control algorithm formulated through Markov Decision Processes (MDPs), including both model-based dynamic programming and optimal control methods, as well as model-free reinforcement learning approaches.
@@ I-B1 Related Works
Information theory provides powerful tools for quantifying the amount of uncertainty in the value of a random variable, and has had a profound influence on the design of communication systems. However, as Howard indicates in his seminal work, “no theory that involves just the probabilities of outcomes without considering their consequences could possibly be adequate in describing the importance of uncertainty to a decision maker” [13]. Howard introduces the theory of VoI, and asserts that numerical values can be assigned to the elimination or reduction of any uncertainty. To explain the concept and intuition of VoI, Howard uses a bidding problem as an example, where the expected profit of eliminating any certainty is used as the VoI. However, Howard’s work does not provide a rigorous definition or mathematical formula for calculating VoI. Following his work, various definitions of VoI have been proposed for different applications and objectives. A survey on VoI in Wireless Sensor Network (WSN) and Internet of Things (IoT) by Alawad et al. [14] provides a structured literature review on the fragmented works in this area. Despite these contributions, a theoretical foundation for VoI is still lacking, which would allow VoI to be estimated using uniform formulas for different applications.
@@ I-B2 Contribution 2
To derive the information theory-based VoI, we calculate the conditional Kullback–Leibler (KL) divergence between the probability distribution given by the product of the transition models of the original state and the V2X information, and the probability distribution given by the transition model of the augmented state, which inclues both the original state and the V2X information. While information theory-based VoI is easier to derive given the model of the environment, utility-based VoI has several important advantages. Therefore, we introduce a DRL-based method to estimate the utility-based VoI by offline training policies with missing/imperfect and full/perfect information, respectively. Then, the regret and advantage function can be estimated to quantify the EVoI and IVoI, respectively.
@@ I-C1 Related Works
In V2X communication, there is a growing interest aimed at designing RRM mechanisms to minimize the Age of Information (AoI) [23, 24, 25, 26], rather than focusing solely on traditional metrics such as throughput and delay. While AoI measures the importance of information by assessing its timeliness, it does not account for the broader benefits that information can provide. In contrast, VoI directly quantifies the advantage of information in enhancing the performance of specific tasks. Unfortunately, there is currently a lack of studies focused on VoI-oriented RRM.
In MDP and RL communities, the problems of “What to Communicate” and “When to Communicate” have been studied in the context of Decentralized-POMDP (Dec-POMDP) [27, 28] and Multi-agent RL (MARL) [29, 30]. The goal in these studies is to simultaneously learn both a communication policy and an acting policy to maximize performance while adhering to communication constraints. The “MARL with communications” framework is expected to play a vital role for 6G use cases involving multiple agents, massive sensory data, and constrained communication resources, especially when the learned communication policies lead to the emergence of agent-to-agent languages [31]. However, VoI is rarely used in this line of research [28]. One negative consequence of this neglect is that it is challenging to explicitly express the trade-off between maximizing the performance of considered tasks and that of traditional communication services. As a result, communication constraints in these studies are often simplified, usually by limiting the size or number of messages, or by restricting which agents are allowed to transmit.
As wireless systems enter the AI era, the semantic communication paradigm, which emphasizes conveying the intended meaning of information rather than individual symbols, has emerged as a promising solution for enhancing wireless performance [32]. Semantic communication and VoI-oriented communication share a common goal-oriented philosophy in communication system design. However, they differ in focus: semantic communication emphasizes understanding the meaning of transmitted content such as text, images, and speech, whereas VoI-oriented communication prioritizes the impact of information on the performance of control tasks supported by communication. Semantic communication is well-suited for non-control services, such as image and text sharing, but may fall short for control-related applications. For instance, in AD, semantic communication focuses on identifying objects like vehicles and roadblocks, while VoI-oriented communication goes further by determining which information is critical for the performance and safety of driving tasks. Therefore, integrating VoI-oriented communication with semantic communication offers a promising direction for supporting control-related services.
@@ II System Model
Figure 1: Illustration of system model.
@@
Figure 2: Examples of typical scenarios.
@@
where $\rho_{i}$ is a time constant representing the driveline dynamics of vehicle $i$ and $u_{i}$ is the vehicle’s control input.
We denote the headway of vehicle $i$, i.e., bumper-to-bumper distance between $i$ and its preceding vehicle (predecessor) $i-1$, by $d_{i}$, which satisfies
@@
$d_{\sigma,i}=\sigma_{i}+h_{i}v_{i},$ (3)
where $\sigma_{i}$ is a constant standstill distance for vehicle $i$ and $h_{i}$ is the desired time-gap of vehicle $i$.
@@
The time horizon is discretized into $K$ control intervals, each with a duration of $T$ seconds. Let $x_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$, where the subscript $k\in\{0,1,\cdots,K-1\}$ indicates the values of the variables at the beginning of control interval $k$. The system dynamics in discrete time are derived from (1) on the basis of forward Euler discretization:
@@
$A_{i}=\begin{bmatrix}1&T&-h_{i}T\\ 0&1&-T\\ 0&0&1-\frac{T}{\rho_{i}}\end{bmatrix},B_{i}=\begin{bmatrix}0\\ 0\\ \frac{T}{\rho_{i}}\end{bmatrix},C_{i}=\begin{bmatrix}0\\ T\\ 0\end{bmatrix}.$ (6)
At the beginning of each control interval, the control agent of vehicle $i$ has to determine the control input $u_{i,k}$ based on the observation of the state. The velocity $v_{i,k}$ and acceleration $acc_{i,k}$ can be measured locally, while the control errors $e_{pi,k}$ and $e_{vi,k}$ can be quantified by a radar unit mounted at the front of the vehicle. On the other hand, vehicle $i$ can only obtain the acceleration $acc_{i-1,k}$ of the predecessor through V2X communications. To determine the optimal action $u_{i,k}$, we must decide when and how to share information by V2X communications from predecessor $i-1$ to its follower $i$.
@@ IV-A SSDP Basics
    which is derived from the predecessor $i-1$’s dynamic model $\dot{acc}_{i-1}=-\frac{1}{\rho_{i-1}}acc_{i-1}+\frac{1}{\rho_{i-1}}u_{i-1}$ based on forward Euler discretization;
@@
    $r(S_{i,k},a_{i,k})=$
    $-\{|\frac{e_{pi,k}}{\hat{e}_{p,\mathrm{max}}}|+a|\frac{e_{vi,k}}{\hat{e}_{v,\mathrm{max}}}|+b|\frac{a_{i,k}}{u_{\mathrm{max}}}|+c|\frac{j_{i,k}}{2acc_{\mathrm{max}}/T}|\},$ (17)
    where $j_{i,k}$ is the change rate in acceleration, known as the jerk:
    $j_{i,k}$ $=\frac{acc_{i,k+1}-acc_{i,k}}{T}=-\frac{1}{\tau_{i}}acc_{i,k}+\frac{1}{\tau_{i}}a_{i,k},$ (18)
    and $\hat{e}_{p,\mathrm{max}}$ and $\hat{e}_{v,\mathrm{max}}$ are the nominal maximum control errors such that it is larger than most possible control errors. $a$, $b$ and $c$ are the positive weights and can be adjusted to determine the relative importance of minimizing each term.
    The reward function reflects the optimization objective: minimizing position error $e_{pi,k}$ and velocity error $e_{vi,k}$ to achieve the platoon control target, while penalizing control input $u_{i,k}$ to reduce fuel consumption and jerk for improved driving comfort.
    To improve the performance of training a DRL agent, a Huber loss function can be used. It applies the squared error instead of the absolute error when the absolute reward value is below a certain threshold [6].
@@ V-A Utility-based VoI
(Value of Missing Information (VoMI)) The VoMI is defined as the difference in performance between the acting policy $\pi$ for the SSDP, where the state $S_{k}$ excludes information $I_{k}$, and the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes $I_{k}$.
@@
(Value of Imperfect Information (VoII)) The VoII is defined as the difference in performance between the acting policy $\tilde{\pi}^{\rm ip}$ for the augmented-state SSDP with imperfect information, where the state $\tilde{S}_{k}^{\mathrm{ip}}$ is constructed from imperfect information $I_{k}^{\rm ip}$, and the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP $\tilde{S}_{k}=\{S_{k},I_{k}\}$, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes perfect information $I_{k}$.
Note that both VoMI and VoII quantify the performance gap between a policy based on partial/imperfect observation (state with partial/imperfect information) and a policy based on full/perfect observation (state with full/perfect information). For ease of discussion, we introduce unified notations $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}$ to denote the policies in the VoMI and VoII definitions based on partial/imperfect and full/perfect observations, respectively - where $\pi^{\rm{inf}}$ yields inferior performance and $\pi^{\rm{sup}}$ exhibits superior performance. Correspondingly, the states are denoted as $S_{k}^{\rm{inf}}$ for partial/imperfect observations and $S_{k}^{\rm{sup}}$ for full/perfect observations.
Taking VoMI as an example, $\pi^{\rm{inf}}$ could be the optimal policy $\pi^{*}$ for the SSDP with the state $S_{k}$ that excludes $I_{k}$. Alternatively, it could be derived from the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP, where $I_{k}$ in $\tilde{S}_{k}=\{S_{k},I_{k}\}$ is replaced with dummy bits. More generally, it may represent any policy employed by the control agent to make decisions based on the partial observation $S_{k}$.
Remark 4.
(Dependence of VoI on policies) In Definitions 5 and 6, we assume that the control agent follows policy $\pi^{\rm{inf}}$ when only partial or imperfect observations are available, and switches to policy $\pi^{\rm{sup}}$ when full or perfect observations are accessible. In other words, the control agent adopts a mixed policy, utilizing $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}$ depending on the quality of the available information, and is assumed to fully exploit the full/perfect information to make more informed decisions. In this setting, the dependence of VoI on the policies is solely determined by the robustness of $\pi^{\rm{inf}}$ to missing or imperfect information: the more robust $\pi^{\rm{inf}}$ is, the less valuable the additional or perfect information becomes.
When $\pi^{\rm{sup}}$ is not available to the control agent, $\pi^{\rm{inf}}$ is used in place of $\pi^{\rm{sup}}$ in both Definitions 5 and 6. In this case, the control agent always employs $\pi^{\rm{inf}}$, regardless of whether it receives partial/imperfect or full/perfect observations. Here, the VoI also depends on how effectively the agent can leverage the additional or perfect information using $\pi^{\rm{inf}}$: the less capable the agent is at utilizing this information, the less valuable it becomes to the agent, even though it may be more useful to another policy that can exploit it more effectively.
@@
Depending on how the performance difference is measured, both VoMI and VoII can be further divided into two categories: Expected Cumulative VoI (EVoI) and Immediate VoI (IVoI).
@@
(Expected Cumulative VoI (EVoI)) The EVoI $\bar{\Xi}$ is defined as the difference in performance between the acting policy $\pi^{\rm{inf}}$ based on partial/imperfect observations and the optimal policy $\pi^{\rm{sup}}$ based on full/perfect observations, respectively. Formally,
@@
The EVoI quantifies the performance loss as a consequence of the control agent consistently receiving the partial/imperfect observations at every control interval. In the terminology of RL, EVoI corresponds to the regret of $\pi^{\rm{inf}}$. It is worth noting that the EVoI is a non-positive value, with a larger EVoI indicating less performance loss. When the EVoI reaches its maximum value of zero, no performance loss occurs, implying that the missing or imperfect information has no impact on decision-making.
@@
(Immediate Value of Information (IVoI)) Given a full/perfect observation $S_{k}^{\rm{sup}}$ and a partial/imperfect observation $S_{k}^{\rm{inf}}$ at control interval $k$, the IVoI $\xi_{k}$ is defined as the performance difference resulting from executing two different actions at interval $k$: one chosen by the acting policy $\pi^{\rm{inf}}$ based on $S_{k}^{\rm{inf}}$ and the other by the optimal policy $\pi^{\rm{sup}}$ based on $S_{k}^{\rm{sup}}$, while following $\pi^{\rm{sup}}$ based on full/perfect observations $S_{k+1}^{\rm{sup}},S_{k+2}^{\rm{sup}},\dots$ thereafter. Formally,
@@
where $A_{\pi^{\rm{sup}}}$ denotes the advantage function under the optimal policy $\pi^{\rm{sup}}$, and the action $a_{k}^{\rm{inf}}$ is sampled according to $\pi^{\rm{inf}}$ based on $S_{k}^{\rm{inf}}$.
The IVoI indicates the performance loss as a consequence of the control agent receiving a partial/imperfect observation only at a specific control interval $k$, while consistently obtaining full/perfect observations throughout the remaining time. It is defined for a given pair consisting of the full/perfect observation $S_{k}^{\rm{sup}}$ and inferior action $a_{k}^{\rm{inf}}$, which is prescribed by the acting policy $\pi^{\rm{inf}}$ based on the partial/imperfect observation $S_{k}^{\rm{inf}}$. In other words, IVoI measures the performance degradation when the agent, despite the full/perfect observation being $S_{k}^{\rm{sup}}=s$, relies on the partial/imperfect observation $S_{k}^{\rm{inf}}$ to select its action at control interval $k$. In the terminology of RL, IVoI corresponds to the advantage function of the optimal policy $\pi^{\rm{sup}}$. Similar to EVoI, IVoI is also a non-positive value, where maximizing IVoI leads to minimizing the performance loss.
@@
Therefore, utility-based VoI can be categorized into two main types: EVoI and IVoI. Each of these types is further divided into two subcategories: VoMI and VoII. As illustrated in Fig. 3, this classification results in four distinct categories of utility-based VoI: Expected VoMI (EVoMI), Expected VoII (EVoII), Immediate VoMI (IVoMI), and Immediate VoII (IVoII). According to Definition 5 and 6, the policy/state/action $\pi^{\rm{inf}}$/$S_{k}^{\rm{inf}}$/$a_{k}^{\rm{inf}}$ in (23) and (24) is replaced by $\pi$/$S_{k}$/$a_{k}$ and $\tilde{\pi}^{\rm ip}$/$\tilde{S}_{k}^{\rm ip}$/$a_{k}^{\rm ip}$ for VoMI and VoII, respectively, while $\pi^{\rm{sup}}$/$S_{k}^{\rm{sup}}$ in (23) and (24) is replaced by $\tilde{\pi}^{*}$/$\tilde{S}_{k}$.
Figure 3: Four categories of Utility-based VoI.